# Magic tricycles: efficient magic state generation with finite block-length quantum LDPC codes

### v1

@@ Appendix E Numerical details of noise simulations
We simulate syndrome extraction circuits for both $X$ and $Z$ checks of the tricycle codes. Each simulation begins by initializing the data qubits in the appropriate basis, followed by multiple cycles of syndrome extraction. At each cycle, fresh ancilla qubits are initialized, entangled with the data via two-qubit gates, and subsequently measured to extract the stabilizer values.
We consider two types of simulations: (1) single-shot simulations in the $Z$ basis, and (2) fault-tolerant $d$-round simulations in the $X$ basis, where the syndrome extraction circuit is repeated for $d$ cycles prior to decoding, with $d$ taken as the upper bound on the code distance.
For both single-shot $Z$-basis and $d$-round $X$-basis simulations, we employ a coloration circuit for syndrome extraction. While we explored various permutations of the entangling gate order within the circuit, only negligible improvements in performance were found. Thus, a random coloration circuit is used in our results. We note that more exhaustive searches over gate orderings could further improve performance, for example by suppressing hook errors.
@@
Decoding is performed using a belief-propagation with ordered statistics decoding (BP+OSD) scheme as presented in Ref. [65]. To enable circuit-level decoding, we map each syndrome extraction circuit to a spacetime code: checks of the spacetime code correspond to parities of stabilizer measurements across consecutive rounds, and the code qubits represent distinct fault locations (e.g., a two-qubit gate error at a specific location in space and time). Formally, for each round, the value of a stabilizer at time $t$ is $m_{t}\in{0,1}$; the detector value is defined as $d_{t}\coloneq m_{t}+m_{t-1}$ (mod $2$). Detectors are associated with the checks of the spacetime code, and the code qubits are assigned to circuit fault locations. A qubit is included in the support of a check if a fault at that location alone would flip the outcome of the associated detector.
@@
Appendix F Optimal-depth syndrome extraction circuits
@@
The above theorem applies to all codes presented in the main text. For other cases, let $W=w_{a}+w_{b}+w_{c}$ and $V=\left|\{w\in\{w_{a},w_{b},w_{c}\}\mid w\text{ is even}\}\right|$. Our scheduling method produces the following result, which is at most two CNOT layers from optimal.
Corollary F.1.1.
@@
Theorem˜F.1 covers $V=0$. For $V=1$, assign the odd-weight polynomial as $\mathbf{A}$, and the previous proof still applies. For $V=2$, assign the odd-weight polynomials as $\mathbf{A}$ and $\mathbf{C}$. Modifying stages 1 and 5 of in Eq.˜40 as follows suffices:
@@
For $V=0$ and $V=1$, initialization precedes and measurement follows the CNOT layers each cycle, adding two units per cycle. For $V=2$ and $V=3$, as illustrated in Fig.˜6a, X-check measurement can be parallelized with the final CNOT layer, and Z-check initialization with the first layer in Eq.˜42, resulting in the stated depths. ∎
@@
Adapt stages 1 and 5 in Eq.˜40 as follows:
@@
where the prime indicates check qubits for the other cycle. This yields a staggered circuit (Fig.˜6b) similar to the construction for hypergraph product codes in Ref. [95]. If at least one polynomial has even weight, assign it as $\mathbf{B}$; if $V=3$, revise stages 2 and 4 similarly to Eq.˜42. ∎
@@
Appendix G Neutral atom array architecture and rearrangement procedure
To concretely specify the atom rearrangement procedure, we consider the reference architecture shown in Fig.˜7, primarily based on settings from recent experiments [4, 5, 23]. The neutral atom array features multiple zones for different functions: an entangling zone for CNOT operations, a storage zone for densely packed atoms, a measurement zone for imaging, and a reservoir for fresh atoms. For our purposes, we focus on the entangling and storage zones.
@@ Appendix G Neutral atom array architecture and rearrangement procedure
The full implementation of the syndrome extraction circuit in Eq.˜40 proceeds similarly. For convenience, the permutation notation is slightly abused: when $i=1$, in expressions like $\mathbf{C}^{T}_{i+w_{c}/2}\mathbf{C}_{i-1+w_{c}/2}[\text{S}_{2},\text{S}_{3},\text{S}_{4}]$, the second matrix refers to the prior permutation on those sectors (e.g., identity for stage (1)), not literally $\mathbf{C}_{w_{c}/2}$. The accompanying video demonstrates an instance with $n_{x}=n_{y}=n_{z}=3$ and $\mathbf{A}=yz^{2}+x^{2}z+x^{2}y+x^{2}y^{2}z$, $\mathbf{B}=yz^{2}+xy+x^{2}y+x^{2}y^{2}z$, and $\mathbf{C}=1+xz^{2}+x^{2}y^{2}+x^{2}y^{2}z$.
@@
We will particularly be interested in the case where $C$ and $D$ are complexes where the cochain spaces are direct sums of $\mathbb{F}_{2}[G]$ itself, as this is the scenario defining group-algebra quantum codes. In this case, the bases for the cochain spaces are given by copies of $G$. For example, we may consider then consider the balanced product of two 22-term co-chain complexes
@@
For module tensor products of $R$ with itself, we have a useful isomorphism $R\otimes_{R}R\cong R$ given by $r\otimes_{R}s\cong rs$ (see Appendix A of [30] for example). This yields the bicycle code chain complex in Eq.˜17. Iterating the balanced product with a third classical complex defined by $c\in\mathbb{F}_{2}[G]$ yields the complex in Eq.˜18.

### v2

@@ Appendix G Numerical details of noise simulations
For both types of simulation, we simulate syndrome extraction circuits for both $X$ and $Z$ checks of the tricycle codes. Each simulation begins by initializing the data qubits in the appropriate basis, followed by multiple cycles of syndrome extraction. At each cycle, fresh ancilla qubits are initialized, entangled with the data via two-qubit gates, and subsequently measured to extract the stabilizer values.
For both single-shot $Z$-basis and $d$-round $X$-basis simulations, we employ a coloration circuit for syndrome extraction together with the standard two-qubit depolarizing circuit-noise error model also used in the main text simulations.
For the single-shot scenario, we utilize a windowed decoding protocol, wherein each window comprises a fixed number of syndrome extraction rounds followed by decoding and correction. This window is then repeated multiple times, providing a concrete probe of the code’s sustainable logical error suppression [20]. Specifically, our simulations employ windows of three rounds, each repeated fourteen times for a total of forty-two rounds—a protocol previously demonstrated to saturate sustainable performance in high-rate codes [107]. With fixed window size, only codes possessing genuine single-shot capability exhibit logical error suppression that scales with code distance; in contrast, codes lacking such features see their performance limited by the window size. Although a window size of one is theoretically sufficient, empirical studies indicate that size three yields enhanced robustness to measurement errors and generally improved logical error rates [107, 14]. For conventional $d$-round simulations in the $X$ basis, the syndrome extraction is repeated for $d$ cycles, and decoding is performed using the entire record of accumulated syndrome information.
@@
Here, we perform decoding using a belief-propagation with ordered statistics decoding (BP+OSD) scheme as presented in Ref. [76]. To enable circuit-level decoding, we map each syndrome extraction circuit to a spacetime code [107, 14]: checks of the spacetime code correspond to parities of stabilizer measurements across consecutive rounds, and the code qubits represent distinct fault locations (e.g., a two-qubit gate error at a specific location in space and time). Formally, for each round, the value of a stabilizer at time $t$ is $m_{t}\in{0,1}$; the detector value is defined as $d_{t}\coloneq m_{t}+m_{t-1}$ (mod $2$). In single-shot simulations, belief propagation is applied to each window, except for the final readout which uses BP+OSD; for $d$-round simulations, BP+OSD is used for decoding the full syndrome record.
@@
We simulate various 4444-4-4 tricycle code and our simulation results are summarized in Fig.˜6. Logical error rate is defined as the probability that any logical qubit within the code block fails under the simulated noise model. We observe a gate-noise threshold greater than $0.4\%$ for each of the $X$ and $Z$ basis results (see the discussion after Theorem˜A.1 in for further details). By fitting the measured logical error rates to an exponential decay model, $p_{L}=\alpha(p/p_{\text{th}})^{\beta n^{\gamma}}$, we estimate the achievable error rates for both $X$ and $Z$ bases. Table˜3 presents logical error rates for two key physical error benchmarks: $p_{2q}=10^{-3}$, representative of near-term hardware, and $10^{-4}$, representative of projected advances, together with relevant code parameters.
For completeness, we also simulate the single-shot performance of the 4224-2-2 tricycle codes explored in the main text. We used the same BP+OSD decoding setup described above. The results can be found in Fig.˜7, which demonstrate impressive single-shot error correction suppression for this class of codes. In practice, however, we anticipate that the overall magic state factory performance will be limited by the $d$-rounds $X$-basis simulated in the main text, owing to its higher check weight and lower distance compared to the $Z$ basis.
Figure 6: Circuit-level noise simulation results for 4444-4-4 tricycle codes. (a) Logical error rate (as measured by block failure probability) versus two-qubit physical gate error rate ($p_{2q}$) for single-shot error correction in the $Z$ basis. Single-shot performance is evaluated using a windowed decoding protocol: three rounds of syndrome extraction followed by decoding and correction, with the window repeated 14 times (for 42 total rounds) to probe sustainable suppression of logical errors. (b) Logical error rate versus $p_{2q}$ for fault-tolerant, $d$-round error correction in the $X$ basis. In both panels, errors are sampled under a standard two-qubit depolarizing circuit-level noise model, and results are shown for various tricycle code. Logical error rates are determined via Monte Carlo simulations using a BP+OSD decoder, with each data point corresponding to $M$ samples; error bars indicate standard errors, computed as $\sqrt{p_{L}(1-p_{L})/M}$.
Table 3: Example 4444-4-4 tricycle codes and their logical error rates under circuit-level noise. Logical error rates $p_{L}^{(X)}$ and $p_{L}^{(Z)}$ are shown for both $p_{2q}=10^{-3}$ and $p_{2q}=10^{-4}$, for different [[N, K, D]] code parameters.
Figure 7: Single-shot simulation results for 4224-2-2 tricycle codes. Logical error rate normalized by number of QEC rounds and number of logical qubits versus two-qubit physical gate error rate ($p_{2q}$) for single-shot error correction in the $Z$ basis. Logical error rates are determined via Monte Carlo simulations using a BP+OSD decoder, with each data point corresponding to $M$ samples; error bars indicate standard errors, computed as $\sqrt{p_{L}(1-p_{L})/M}$.
Appendix H Optimal-depth syndrome extraction circuits
@@
The construction allows certain degrees of freedom. First, the assignment of the matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ to stages 1+5, 2+4, and 3 can be chosen arbitrarily. Second, the ordering of layers within each stage is flexible, provided that stages 1 and 5 contain complementary terms, as do stages 2 and 4. Let $w_{1}$, $w_{2}$, and $w_{3}$ denote the weights of polynomials assigned to stages 1+5, 2+4, and 3, respectively. (In the proof above, $w_{1}=w_{c}$, $w_{2}=w_{b}$, and $w_{3}=w_{a}$.) Then, the layers within stage 3 can be ordered $w_{3}!$ ways. The order of terms in stages 1 and 5 for the X-checks can be arbitrary, resulting in $w_{1}!$ possibilities. Since stages 1 and 5 must be complementary for the Z-checks, the assignment of terms to stages 1 and 5 is fixed, but permutations within each stage remain possible, contributing an additional $((w_{1}/2)!)^{2}$ arrangements. Similarly, for stages 2 and 4, the same counting applies with weight $w_{2}$. Thus, the total number of distinct circuit configurations is
$w_{1}!\times\bigl((w_{1}/2)!\bigr)^{2}\times w_{2}!\times\bigl((w_{2}/2)!\bigr)^{2}\times w_{3}!.$ (71)
For example, when $w_{1}=w_{2}=2$ and $w_{3}=4$, there are $96$ circuits, while for $w_{1}=w_{3}=2$ and $w_{2}=4$, there are $384$ circuits. In addition, the assignments of $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ to the stages 1+5, 2+4, and 3 themselves can be permuted. For the 4224-2-2 codes, this yields: two assignments with $w_{1}=w_{2}=2,w_{3}=4$, two assignments with $w_{1}=w_{3}=2,w_{2}=4$, and two assignments with $w_{2}=w_{3}=2,w_{1}=4$. Altogether, this gives $96\times 2+384\times 4=$ 1,728 different circuits for the 4224-2-2 codes. A similar calculation for the 4424-4-2 and 4444-4-4 codes results in 55,296 and 1,327,104 distinct circuits, respectively.
In the counting above, same polynomial terms are scheduled together if possible, to enable more efficient atom rearrangement (more details in Appendix˜I). For example, in stage 3, $\mathbf{A}_{1}$ from $X$ to $D_{1}$, from $D_{2}$ to $Z_{3}$, and from $D_{3}$ to $Z_{1}$ are always in the same layer, but they do not have to be. If we consider such possibilities, the number of different choices for stage 3 is no longer $w_{3}!$ but $(w_{3}!)^{3}$. For a specific choice of the weights, the number of different circuits becomes
$w_{1}!\times\bigl((w_{1}/2)!\bigr)^{4}\times w_{2}!\times\bigl((w_{2}/2)!\bigr)^{4}\times(w_{3}!)^{3}.$ (72)
For 4224-2-2 codes, this amounts to 135,168 circuits.
Theorem˜H.1 applies to all codes presented in the main text. For other cases, let $W=w_{a}+w_{b}+w_{c}$ and $V=\left|\{w\in\{w_{a},w_{b},w_{c}\}\mid w\text{ is even}\}\right|$. Our scheduling method produces the following result, which is at most two CNOT layers from optimal.
Corollary H.1.1.
@@
Theorem˜H.1 covers $V=0$. For $V=1$, assign the odd-weight polynomial as $\mathbf{A}$, and the previous proof still applies. For $V=2$, assign the odd-weight polynomials as $\mathbf{A}$ and $\mathbf{C}$. Modifying stages 1 and 5 of in Eq.˜69 as follows suffices:
@@
For $V=0$ and $V=1$, initialization precedes and measurement follows the CNOT layers each cycle, adding two units per cycle. For $V=2$ and $V=3$, as illustrated in Fig.˜8a, X-check measurement can be parallelized with the final CNOT layer, and Z-check initialization with the first layer in Eq.˜73, resulting in the stated depths. ∎
@@
Adapt stages 1 and 5 in Eq.˜69 as follows:
@@
where the prime indicates check qubits for the other cycle. This yields a staggered circuit (Fig.˜8b) similar to the construction for hypergraph product codes in Ref. [107]. If at least one polynomial has even weight, assign it as $\mathbf{B}$; if $V=3$, revise stages 2 and 4 similarly to Eq.˜73. ∎
@@
Appendix I Neutral atom array architecture and rearrangement procedure
To concretely specify the atom rearrangement procedure, we consider the reference architecture shown in Fig.˜9, primarily based on settings from recent experiments [6, 7, 25]. The neutral atom array features multiple zones for different functions: an entangling zone for CNOT operations, a storage zone for densely packed atoms, a measurement zone for imaging, and a reservoir for fresh atoms. For our purposes, we focus on the entangling and storage zones.
@@ Appendix I Neutral atom array architecture and rearrangement procedure
The full implementation of the syndrome extraction circuit in Eq.˜69 proceeds similarly. For convenience, the permutation notation is slightly abused: when $i=1$, in expressions like $\mathbf{C}^{T}_{i+w_{c}/2}\mathbf{C}_{i-1+w_{c}/2}[\text{S}_{2},\text{S}_{3},\text{S}_{4}]$, the second matrix refers to the prior permutation on those sectors (e.g., identity for stage (1)), not literally $\mathbf{C}_{w_{c}/2}$. The accompanying video demonstrates an instance with $n_{x}=n_{y}=n_{z}=3$ and $\mathbf{A}=yz^{2}+x^{2}z+x^{2}y+x^{2}y^{2}z$, $\mathbf{B}=yz^{2}+xy+x^{2}y+x^{2}y^{2}z$, and $\mathbf{C}=1+xz^{2}+x^{2}y^{2}+x^{2}y^{2}z$.
@@
We will particularly be interested in the case where $C$ and $D$ are complexes where the cochain spaces are direct sums of $\mathbb{F}_{2}[G]$ itself, as this is the scenario defining group-algebra quantum codes. In this case, the bases for the cochain spaces are given by copies of $G$. For example, we may consider then consider the balanced product of two 22-term cochain complexes
@@
For module tensor products of $R$ with itself, we have a useful isomorphism $R\otimes_{R}R\cong R$ given by $r\otimes_{R}s\cong rs$ (see Appendix A of [33] for example). This yields the bicycle code chain complex in Eq.˜24. Iterating the balanced product with a third classical complex defined by $c\in\mathbb{F}_{2}[G]$ yields the complex in Eq.˜25.